OpenStudy (anonymous):

6.03 Vectors in the Plane 1.Let u = <-4, -3>. Find the unit vector in the direction of u, and write your answer in component form. 2.Let u = <-8, -9>. Find 7u Confused, please help!

OpenStudy (sirm3d):

the unit vector in the direction of u is \[\large \frac{1}{|u|}u\]

OpenStudy (sirm3d):

if k is scalar, k<a,b> = <ka. kb>

OpenStudy (sirm3d):

if u=<a,b> then |u| = sqrt(a^2 + b^2)

OpenStudy (anonymous):

The first equation you posted that would be to solve # 1 right? I would convert the vector to slope and then plug that into U?

OpenStudy (anonymous):

Oh, wait I see, I would use sqrt(a^2 + b^2) then plug it in to U. But what about the U that isn't absolute value?

OpenStudy (sirm3d):

the u that isn't inside the absolute value signs is your given vector

OpenStudy (anonymous):

My given vector is <-4, -3> correct?

OpenStudy (sirm3d):

yes. you want the unit vector in the same direction as that given vector

OpenStudy (sirm3d):

that should be sqrt(25), not 25, in the denominator

OpenStudy (anonymous):

1/ sqrt(-4^2 + -3^2) 1/5 (u)

OpenStudy (sirm3d):

just "distribute" the scalar into the components of the vector, like k<a,b> = <ka, kb>

OpenStudy (anonymous):

Oh, Okay, So: 1/5(-4) = -.8 Or -4/5 1/5(-3) = -.6 Or -3/5

OpenStudy (sirm3d):

yes.

OpenStudy (anonymous):

Thank you so much!! I would do the same for #2 but multiply by 7 right?

OpenStudy (anonymous):

answer for # 2 = <-56, -63> thanks again

OpenStudy (sirm3d):

right.